OpenStudy (aundiewear):

sin(105) = the answer is (sqrt(2) + sqrt(6)) / 4 but I don't get how to get how

OpenStudy (anonymous):

\[105=60+45\]

OpenStudy (anonymous):

use the "addition angle" formula for sine

OpenStudy (anonymous):

\[\sin(A+B)=\sin(A)\cos(B)+\sin(B)\cos(A)\] replace \(A\) by \(60^\circ\) and \(B\) by \(45^\circ\)

OpenStudy (anonymous):

alright alright get this dude, the answer might be more then 12

OpenStudy (anonymous):

that would be a miracle since sine is never larger than one

OpenStudy (anonymous):

i said might come on dudes i look @ signs on the road all the time